# Revolutionizing Command Interface: Maximizing Control Efficiency in INO ICAL Experiment with UDP Protocol

### v1

@@ 7 Other Responsibilities of Command server in ICAL Experiment
HPCI allows the command server to configure run conditions including setting up TCP sockets for event acquisition, enabling periodic monitoring, and facilitating remote system upgrades. These processes are essential for seamless data collection and system maintenance. HPCI is instrumental in distributing epoch time to all the DAQs. This timestamp is used by the Real Time Clock (RTC) hardware logic in RPC-DAQ to precisely synchronize commands and events. This helps in accurate event time-stamping and data correlation by the event servers. The Detector parameters like high voltage (HV) and channel masking are configured and monitored through HPCI. This allows the detectors to operate with specified parameters and maintain optimal condition. HPCI allows user to configure network parameters online. These settings includes IP addresses, Multicasting addresses, port numbers, and socket mode settings (TCP/UDP). HPCI provides a centralized interface for seamless network configuration adjustments.
@@ 8 Conclusion
The HPCI approach is not a novel concept in network-based DAQs, its application within the INO ICAL project presents unique challenges due to the sheer scale of network-enabled DAQs involved. So HPCI is devised with traditional command interface approach flavored with smarter handshaking techniques. Currently, HPCI based command server is scaled and tested upto 20 RPC-DAQs in mICAL. However, the real test lies ahead as it will soon be deployed and evaluated with a significantly larger number of DAQs within the INO ICAL project. The HPCI continues to evolve and time tested in Mini-ICAL.

### v2

@@ 8 Other Responsibilities of Command server in ICAL Experiment
HPCI allows the command server to configure run conditions including setting up TCP sockets for event acquisition, enabling periodic monitoring, and facilitating remote system upgrades. These processes are essential for seamless data collection and system maintenance. HPCI is instrumental in distributing epoch time to all the DAQs. This timestamp is used by the Real Time Clock (RTC) hardware logic in RPC-DAQ to precisely synchronize commands and events. This helps in accurate event time-stamping and data correlation by the event servers. The Detector parameters like high voltage (HV) and channel masking are configured and monitored through HPCI. This allows the detectors to operate with specified parameters and maintain optimal condition. Due to the large number of RPC-DAQ modules in the INO ICAL experiment, a scalable configuration management is needed. Each DAQ is equipped with programmable flash memory that stores critical network parameters including IPv4 address, MAC address, subnet mask, gateway and multicast group information. These settings are read by the embedded soft-core processor during power-up, ensuring the module initializes with correct network settings. To handle remote management, HPCI Back-end supports a dedicated set of configuration commands that allow the central server to modify these parameters dynamically. This configuration mechanism is integrated within the run-control software, which acts as a centralized utility for managing DAQ identity and network mapping. The DAQ configuration can be updated over the network and reloaded during controlled reboot cycles. While general-purpose configuration tools like Puppet or Ansible are typically used for server clusters, our approach reflects the practical constraints of embedded systems without operating systems. A similar mechanism is used in FPGA-based systems such as IPBus [6] and the RFU command protocol [14], where networked control and addressability are managed directly through firmware-level command stacks.
@@ 9 Conclusion
Although the HPCI approach is not a novel concept in network-based DAQs, its application within the INO ICAL project presents unique challenges due to the sheer scale of network-enabled DAQs involved. As a consequence, HPCI is devised with traditional command interface approach flavored with smarter handshaking techniques. Currently, HPCI based command server is scaled and tested up to 20 RPC-DAQs in mini-ICAL. However, the real test lies ahead as it will soon be deployed and evaluated with a significantly larger number of DAQs within the INO ICAL project. The HPCI continues to evolve and be time tested in mini-ICAL.